Systems Staff Meeting Notes 6/12/00

Systems Staff Meeting Notes
June 12, 2000

Present: Rich Bennett, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Suzy Shaw, Hilton Soto, Ying Tang
Announcements
  1. Lindsay's replacement, Andrew Zahn, has been hired. They will work in tandem until her last day, June 23.
  2. A fire suppression system inspection has been scheduled for June 19. Hopefully this time they will not find a leak in the tank.
  3. Bill Covey will be out on Professional Development Leave for the entire fall semester. He will be getting up to speed on NT migration issues. Martha Hruska will be in charge during his departure.
Review Minutes of 6/5/00
  1. Everett's temporary 6-month position will be advertised starting today.
  2. DLC software coordinator interviews are complete. Very good crop of candidates, decision will be tough.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. 10 work stations, 5 laptops are in the queue.
  2. There will probably be one more order before the end of the fiscal year.
Liaisons
  1. Everett & Diana delivered their liaison training committee report to the liaisons.
Networking
  1. Switch problems - Will had to reinstall the software on a switch in MSL last Friday. It wouldn't communicate, he made sure the spanning tree was turned off & it lost power. When rebooted, he discovered that it resets itself to turn the spanning tree on. You have to save the configuration or make the hyperterminal command to turn off the spanning tree. This helps to explain some problems with switches that we've had in the past.
  2. We recently had a case of a person who was unable to send mail. Turns out, the machine was completely infected and the virus checker would not let mail go out. The incoming mail from the infected station was getting recorded in the Trend Scan Mail log.
  3. New Outlook patch from Microsoft - will eliminate any possibility of a Melissa-type virus, but it is not readily removable & would gut many of the features currently in use with Outlook. We're going to hold off on advocating installing the patch in the hopes that they will come out with something in between the severity of this patch. If not, we may eventually have to install it.
  4. NT9 - has been stable for a week, will move more print queues back onto it. Currently, only a test printer & one in DLC are on it.
  5. West - should be finished today. Once completed, we'll have to do some rerouting.
  6. Novell CD Conversion - need to know a good time to switch out servers - probably early AM or on the weekend. During the intercession would be even better - Summer A ends on Friday, June 23. We'll wait until then to proceed. The public work stations need to have access, will have to advertise it as Smathers.
  7. Wireless - still waiting on conduit, Brad's priority list keeps growing.
  8. Documentation list of switches & hubs - West list is current, there is only one switch located in Documents.
Software Projects Status
  1. Preservation Registry migration - this project is ongoing, it's the software that Special Collections uses to send brittle books for repair to John Freund.
  2. Q&A - Hilton & Winston met with Gus, Nelda & Cathy - migrated the test data. They want to test it for a few weeks; they have some data cleanup to do before the real copy is migrated.
  3. Personnel - system modifications are complete. Cube is installed and operating, Winston will meet with Melanie regarding one last thing.
  4. Library Statistics - Cecilia is able to get to the input mechanism, but she is having some other computer problems.
  5. LASSIE - Hilton made the changes requested by Resource Services. So far, so good.
  6. IR Database Cube - the systems tech staff still needs to get together with Ying for training.
  7. Systems wiring - have recently been discussing other ways to display the information. Ying is looking into Visio software - to integrate it for a database view to create diagrams.
DP Coordinator
  1. Students are resuming full-scale tape cleaning.
  2. Debra is copying old Census tapes; Bill anticipates heavy usage by users doing cross group comparisons.
  3. Had an inquiry from NERDC asking if Censpac was still useable. Looks like we may have the only copies of some of this data.
  4. It is unknown if we will be receiving any 2000 Census data in tape format. May just be released optically, or it may go to a third party for conversion.
Calendar
  1. Everett is out today & June 16; then he's gone for 6 months. We'll do lunch with him on Thursday.
  2. Winston out June 19-23.
  3. Bobbie out July 10-21.
  4. Will out July 24-28.


Back to the Systems Homepage

Last updated June 19, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu